The volume of a right cylinder is 144pi cubic inches and it's height is 4 inches. Find the the total area of the solid.

Answer by amarjeeth123(574)

Volume of a right cylinder=π*r^2*h where r is the radius and h is the height
The volume of a right cylinder is 144pi cubic inches and it's height is 4 inches.
Plugging in the values we get π*r^2*4=144*π
r^2=36
r=6 inches
Total area of the solid=2*π*r(r+h)=2*π*6(6+4)=120π square inches.